Aki's Bicycle Designs has determined that when x hundred bicycles are​ built, the average cost per bicycle is given by​ C(x)equals=0.20.2xsquared2minus−1.11.1xplus+9.7399.739​, where​ C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per​ bicycle?
Its C(x)=0.2x(squared)-1.1X+9.739 . . .

Answer by rothauserc(4718)

The graph of C(x) is a parabola that curves upward, we want the x coordinate of the parabola's vertex
:
C(x) = 0.2x^2 -1.1x +9.739
:
For C(x), a = 0.2, b = -1.1
:
the x coordinate of the vertex is -b / 2a = -(-1.1) / 2(0.2) = 2.75
:
we substitute 2.75 for x in the function C(x) to find minimum cost
:
C(2.75) = 0.2(2.75)^2 -1.1(2.75) + 9.739 = 8.2265
:
***********************************************************
2.75 * 100 = 275
Aki Bicycle Designs can build 275 bicycles at minimum cost
***********************************************************
: